# Sums related to Euler’s totient function

### v6

@@ 1. Introduction
[…]
Let $\alpha\in(0,1]$, $y=\max((\ln M)^{\alpha},2)$, and $\mathcal{D}$ be the collection of square-free numbers, all of whose prime divisors lie in $(1,y]$ (we note that $1\in\mathcal{D}$). Let
$\omega(n)\leq\frac{K}{g(n)}$
@@
$\sum_{n=1}^{N}\bigg{(}\frac{a_{n}}{\varphi(a_{n})}\bigg{)}^{s}\leq K\Big{(}\frac{c}{\alpha}\Big{)}^{s}\prod_{p\leq y}\Big{(}1+\frac{(1+p^{-1})^{s}-1}{g(p)}\Big{)}$
@@
Suppose that the assumptions of Theorem 1.1 hold, $K=\gamma N$, $g(p)\geq c_{0}>0$ for any prime $p$, and
$L=\sum_{p}\frac{1}{pg(p)}<\infty.$
@@
#$\#\big{\{}n\leq N:\frac{a_{n}}{\varphi(a_{n})}>t\big{\}}\leq c_{1}\textup{exp}(-\textup{exp}(c_{2}t))N$ (1.2)
@@
#$\#\big{\{}n\leq x:\frac{f(n)}{\varphi(f(n))}>t\big{\}}\leq c_{1}\textup{exp}(-\textup{exp}(c_{2}t))x.$ (1.4)
@@
Let =$\mathcal{L}=\{L_{1},\ldots,L_{k}\}$ be a set of $k$ distinct linear functions with integer coefficients
@@
$\sum_{\begin{subarray}{c}|b|\leq\eta\log x\\ b\,\neq\,b_{1},\ldots,b_{k}\end{subarray}}\Big{(}\frac{a^{k+1}|f(b)|}{\varphi(a^{k+1}|f(b)|)}\Big{)}^{s}\leq\Big{(}c\,\frac{a}{\varphi(a)}\log k\Big{)}^{s}\eta\log x.$ (1.5)
@@
$\sum_{\begin{subarray}{c}|b|\leq\eta\log x\\ b\,\neq\,b_{1},\ldots,b_{k}\end{subarray}}\Big{(}\frac{a^{k+1}|f(b)|}{\varphi(a^{k+1}|f(b)|)}\Big{)}^{s}\leq\Big{(}c\,\frac{a}{\varphi(a)}\log s\Big{)}^{s}\eta\log x.$ (1.6)
@@
Theorem 1.4 extends a result of Maynard ([2, Lemma 8.1]) which showed the same result but with $s=1$ and extends Theorem 1.4 in [3] which showed the inequality (1.6) but with an upper bound
@@
#$\#\big{\{}p\leq x:\frac{f(p)}{\varphi(f(p))}>t\big{\}}\leq c_{1}\textup{exp}(-\textup{exp}(c_{2}t))\pi(x).$ (1.8)
In particular, we can take $f(x)=x-1$ or $f(x)=x^{2}+1$ in Theorem 1.5.
@@ 2. Notation
We reserve the letter $p$ for primes. In particular, the sum $\sum_{p\leq K}$ should be interpreted as being over all prime numbers not exceeding $K$. We denote the number of primes not exceeding $x$ by $\pi(x)$ and the number of primes not exceeding $x$ that are congruent to $a$ modulo $k$ by $\pi(x;k,a)$. Let #$\#A$ denote the number of elements of a finite set $A$. We write $\mathbb{Z}$ for the set of all integers, $\mathbb{N}$ for the set of all positive integers, $\mathbb{R}$ for the set of all real numbers, and $\mathbb{P}$ for the set of all primes. Let $(a_{1},\ldots,a_{n})$ be the greatest common divisor of integers $a_{1},\ldots,a_{n}$ and $[a_{1},\ldots,a_{n}]$ their least common multiple. For real numbers $x,$ $y$ we also use $(x,y)$ to denote the open interval and $[x,y]$ to denote the closed interval. The usage of the notation will be clear from the context.
@@
$\frac{n}{\varphi(n)}\leq\frac{c}{\alpha}\prod_{p|n:\,\,p\leq(\ln n)^{\alpha}}\biggl{(}1+\frac{1}{p}\biggr{)}.$
@@
$\frac{a_{n}}{\varphi(a_{n})}$ $\leq\frac{c}{\alpha}\prod_{p|a_{n}:\,\,p\leq(\ln a_{n})^{\alpha}}\biggl{(}1+\frac{1}{p}\biggr{)}$
@@
$\leq\Big{(}\frac{c}{\alpha}\Big{)}^{s}\sum_{d_{1},\ldots,d_{s}\in\mathcal{D}}\frac{\mu^{2}(d_{1})\ldots\mu^{2}(d_{s})}{d_{1}\ldots d_{s}}\sum_{\begin{subarray}{c}1\leq n\leq N:\\ d_{1}|a_{n},\ldots,\,d_{s}|a_{n}\end{subarray}}1$
@@
It is easy to see that $f(n)$ is supported on $\mathcal{D}$. Suppose that $m,n\in\mathcal{D}$ and $(m,n)=1$. For any $d_{1},\ldots,d_{s}\in\mathcal{D}$ such that $[d_{1},\ldots,d_{s}]=mn$ there are unique $d_{i}^{{}^{\prime}},d_{i}^{{}^{\prime\prime}}\in\mathcal{D},$ $i=1,\ldots,s,$ such that $d_{i}^{{}^{\prime}}d_{i}^{{}^{\prime\prime}}=d_{i}$, $(d_{i}^{{}^{\prime}},d_{i}^{{}^{\prime\prime}})=1$, $[d_{1}^{{}^{\prime}},\ldots,d_{s}^{{}^{\prime}}]=n$, and $[d_{1}^{{}^{\prime\prime}},\ldots,d_{s}^{{}^{\prime\prime}}]=m$. Hence
@@
Since $\omega(n)\leq K/g(n)$ for any $n\in\mathcal{D}$, we have
$\sum_{n=1}^{N}\Big{(}\frac{a_{n}}{\varphi(a_{n})}\Big{)}^{s}$ $\leq K\Big{(}\frac{c}{\alpha}\Big{)}^{s}\sum_{n\in\mathcal{D}}\frac{f(n)}{g(n)}=K\Big{(}\frac{c}{\alpha}\Big{)}^{s}\prod_{p\leq y}\Big{(}1+\frac{f(p)}{g(p)}\Big{)}$
$=K\Big{(}\frac{c}{\alpha}\Big{)}^{s}\prod_{p\leq y}\Big{(}1+\frac{(1+p^{-1})^{s}-1}{g(p)}\Big{)}.$
@@
$\sum_{n=1}^{N}\bigg{(}\frac{a_{n}}{\varphi(a_{n})}\bigg{)}^{s}$ $\leq\gamma N\Big{(}\frac{c}{\alpha}\Big{)}^{s}\prod_{p\leq y}\Big{(}1+\frac{(1+p^{-1})^{s}-1}{g(p)}\Big{)}$
$\leq\gamma N\Big{(}\frac{c}{\alpha}\Big{)}^{s}\prod_{p}\Big{(}1+\frac{(1+p^{-1})^{s}-1}{g(p)}\Big{)}.$ (3.1)
@@
$\prod_{p>s}\Big{(}1+\frac{(1+p^{-1})^{s}-1}{g(p)}\Big{)}\leq\prod_{p>s}\Big{(}1+\frac{es}{pg(p)}\Big{)}\leq\textup{exp}\Big{(}es\sum_{p>s}\frac{1}{pg(p)}\Big{)}\leq\textup{exp}(eLs).$ (3.3)
@@
$1+\frac{(1+p^{-1})^{s}-1}{g(p)}<1+\frac{(1+p^{-1})^{s}}{c_{0}}\leq\Big{(}1+\frac{1}{c_{0}}\Big{)}(1+p^{-1})^{s}\leq\Big{(}1+\frac{1}{c_{0}}\Big{)}\textup{exp}\Big{(}\frac{s}{p}\Big{)}.$
@@
$\prod_{p\leq s}\Big{(}1+\frac{(1+p^{-1})^{s}-1}{g(p)}\Big{)}$ $\leq\Big{(}1+\frac{1}{c_{0}}\Big{)}^{s}\textup{exp}\Big{(}\sum_{p\leq s}\frac{s}{p}\Big{)}$
$\leq\Big{(}1+\frac{1}{c_{0}}\Big{)}^{s}\textup{exp}(s\log\log(s+2)+c_{3}s),$ (3.4)
where $c_{3}>0$ is an absolute constant. From (3.1), (3.3), and (3.4) we obtain
@@
#$\#\Big{\{}n\leq N:\frac{a_{n}}{\varphi(a_{n})}>t\Big{\}}\leq\textup{exp}(s\log\log(s+2)+Cs-s\log t)N.$ (3.5)
@@
#$\#\Big{\{}n\leq N:\frac{a_{n}}{\varphi(a_{n})}>t\Big{\}}\leq\textup{exp}(-\textup{exp}(c_{2}t))N,$
@@
#$\#\Big{\{}n\leq N:\frac{a_{n}}{\varphi(a_{n})}>t\Big{\}}$ $\leq N\leq N\,\frac{\textup{exp}(\textup{exp}(c_{2}t_{0}))}{\textup{exp}(\textup{exp}(c_{2}t))}$
@@
#$\#\Big{\{}n\leq N:\frac{a_{n}}{\varphi(a_{n})}>t\Big{\}}\leq c_{1}\textup{exp}(-\textup{exp}(c_{2}t))N$
@@
where $\gamma=2cd$. The function $g(n)=n^{1/d}$ is multiplicative. Also, $g(p)=p^{1/d}\geq 1$ for any prime $p$ (hence, we can take $c_{0}=1$) and
$L=\sum_{p}\frac{1}{pg(p)}=\sum_{p}\frac{1}{p^{1+1/d}}<\infty.$
@@
#$\#\Big{\{}n\leq x:\frac{f(n)}{\varphi(f(n))}>t\Big{\}}\leq\frac{\widetilde{c}_{1}}{\textup{exp}(\textup{exp}(\widetilde{c}_{2}t))}x.$
@@
$S=\sum_{b\in\Omega}\Big{(}\frac{a^{k+1}|f(b)|}{\varphi(a^{k+1}|f(b)|)}\Big{)}^{s}\leq\Big{(}\frac{a}{\varphi(a)}\Big{)}^{s}\sum_{b\in\Omega}\Big{(}\frac{|f(b)|}{\varphi(|f(b)|)}\Big{)}^{s}.$ (4.3)
@@
$g(p^{\beta})=\begin{cases}p/\min(p,k),&\text{if $\beta=1$;}\\ 1,&\text{if $\beta\geq 2$};\end{cases}$
@@
$\sum_{b\in\Omega}\Big{(}\frac{|f(b)|}{\varphi(|f(b)|)}\Big{)}^{s}\leq c^{s}(\eta\log x)\prod_{p\leq y}\Big{(}1+\frac{(1+p^{-1})^{s}-1}{g(p)}\Big{)}$
$\leq c^{s}(\eta\log x)\prod_{p}\Big{(}1+\frac{(1+p^{-1})^{s}-1}{g(p)}\Big{)}$ (4.4)
@@
$S\leq\Big{(}c\,\frac{a}{\varphi(a)}\Big{)}^{s}\eta\log x\prod_{p}\Big{(}1+\frac{(1+p^{-1})^{s}-1}{g(p)}\Big{)}.$
Suppose that $s\leq k$ (we recall that $2\leq k<\log\log x$). If $p>k$, then $p>s$ and (3.2) holds. Since $g(p)=p/k$, we obtain
$\prod_{p>k}\Big{(}1+\frac{(1+p^{-1})^{s}-1}{g(p)}\Big{)}\leq\prod_{p>k}\Big{(}1+\frac{eks}{p^{2}}\Big{)}\leq c_{1}^{s},$
@@
$\prod_{p\leq k}\Big{(}1+\frac{(1+p^{-1})^{s}-1}{g(p)}\Big{)}\leq\textup{exp}\Big{(}\sum_{p\leq k}\frac{s}{p}\Big{)}\leq(c_{2}\log k)^{s},$
@@
Suppose that $s>k$ and $2\leq k<\log\log x$. Since $g(p)\geq p/k>p/s$ for any prime $p$, we obtain
$\prod_{p>s}\Big{(}1+\frac{(1+p^{-1})^{s}-1}{g(p)}\Big{)}\leq\prod_{p>s}\Big{(}1+\frac{es^{2}}{p^{2}}\Big{)}\leq c_{3}^{s}.$
Since $g(p)\geq 1$ for any prime $p$, we have
$\prod_{p\leq s}\Big{(}1+\frac{(1+p^{-1})^{s}-1}{g(p)}\Big{)}\leq\textup{exp}\Big{(}\sum_{p\leq s}\frac{s}{p}\Big{)}\leq(c_{4}\log s)^{s}.$
@@
We set $a_{p}=f(p)$. It is clear that $N=\pi(x)$. There exists a positive constant $\tau$ depending only on $f$ such that $f(n)\leq\tau n^{d}$ for any positive integer $n$. We take $M=\tau x^{d}$ and $\alpha=1/2$. We assume that $x\geq x_{0}$, where $x_{0}>0$ is a large constant depending only on $f$. By (4.1) and (4.2) we have $y=(\log M)^{\alpha}\leq(\log x)^{3/4}$ and $n\leq\sqrt{x}$ for any $n\in\mathcal{D}$.
We need the following result (see, for example, [4, Exercise 9.3.13]).
@@
For any $\varepsilon>0$ there exists a constant $x_{0}(\varepsilon)>0$ depending only on $\varepsilon$ such that for any $a\in\mathbb{Z}$, $k\in\mathbb{N}$, and $x\in\mathbb{R}$ with $(a,k)=1$, $k\leq x$ and $x\geq x_{0}(\varepsilon)$ we have
$\pi(x;k,a)\leq\frac{(2+\varepsilon)x}{\varphi(k)\log(2x/k)}.$
Fix $n\in\mathcal{D}$, and let $\alpha_{1},\ldots,\alpha_{l}$ be all solutions of the congruence $f(x)\equiv 0$ (mod $n$). For any prime $p|n$, by Lagrange’s theorem we have $\rho(f,p)\leq\min(p,d)$. Hence,
$l=\rho(f,n)=\prod_{p|n}\rho(f,p)\leq\prod_{p|n}\min(p,d).$ (5.1)
@@
Let $1\leq i\leq s$, and suppose that $(\alpha_{i},n)=1$. We have
$\log\Big{(}\frac{2x}{n}\Big{)}\geq\log(2\sqrt{x})\geq\frac{1}{2}\log x.$
Applying Lemma 5.1 with $\varepsilon=1/2$, we obtain
$\pi(x;n,\alpha_{i})\leq\frac{5x}{\varphi(n)\log x},$ (5.3)
if $x_{0}$ is large enough.
@@
$\pi(x;n,\alpha_{i})\leq 1\leq\frac{5x}{\varphi(n)\log x},$
@@
$\frac{5x}{\varphi(n)\log x}\geq\frac{5x}{n\log x}\geq\frac{5\sqrt{x}}{\log x}>1,$
if $x_{0}$ is large enough. We obtain that the inequality (5.3) holds in both cases $(\alpha_{i},n)=1$ and $(\alpha_{i},n)>1$. From (5.1) – (5.3) we get
$\omega(n)\leq\frac{5xl}{\varphi(n)\log x}\leq\frac{\gamma\pi(x)}{\varphi(n)}\prod_{p|n}\min(p,d),$
where $\gamma>0$ is an absolute constant.
@@
$g(p^{\beta})=\begin{cases}(p-1)/\min(p,d),&\text{if $\beta=1$;}\\ 1,&\text{if $\beta\geq 2$};\end{cases}$
@@
$L=\sum_{p}\frac{1}{pg(p)}=\sum_{p\leq d}\frac{1}{p-1}+\sum_{p>d}\frac{d}{p(p-1)}<\infty,$
by Theorem 1.2 there exist positive constants $C$, $c_{1}$ and $c_{2}$ depending only on $d$ such that the inequalities (1.7) and (1.8) hold.
@@
#$\#\Big{\{}p\leq x:\frac{f(p)}{\varphi(f(p))}>t\Big{\}}\leq\frac{\widetilde{c}_{1}}{\textup{exp}(\textup{exp}(\widetilde{c}_{2}t))}\pi(x).$

### v7

@@ 1. Introduction
$\omega(d)=\#\{n\leq N:a_{n}\equiv 0\text{ {(mod $d$)}}\}.$
Let $\alpha\in(0,1]$, $y=\max((\log M)^{\alpha},2)$, and $\mathcal{D}$ be the collection of square-free numbers, all of whose prime divisors lie in $(1,y]$ (we note that $1\in\mathcal{D}$). Let
$\omega(n)\leq Kg(n)$
@@
$\sum_{n=1}^{N}\bigg(\frac{a_{n}}{\varphi(a_{n})}\bigg)^{s}\leq K\Big(\frac{c}{\alpha}\Big)^{s}\prod_{p\leq y}\Big(1+((1+p^{-1})^{s}-1)g(p)\Big)$
@@
Suppose that the assumptions of Theorem 1.1 hold, $K=\gamma N$, $0<g(p)\leq c_{0}$ for any prime $p$, and
$L=\sum_{p}\frac{g(p)}{p}<\infty.$
@@
$\Big\{n\leq N:\frac{a_{n}}{\varphi(a_{n})}>t\Big\}\leq c_{1}\textup{exp}(-\textup{exp}(c_{2}t))N$ (1.2)
@@
$\Big\{n\leq x:\frac{f(n)}{\varphi(f(n))}>t\Big\}\leq c_{1}\textup{exp}(-\textup{exp}(c_{2}t))x.$ (1.4)
@@
Let $\mathcal{L}=\{L_{1},\ldots,L_{k}\}$ be a set of $k$ distinct linear functions with integer coefficients
@@
$\sum_{\begin{subarray}{c}|b|\leq\eta\log x\\ b\,\neq\,b_{1},\ldots,b_{k}\end{subarray}}\bigg(\frac{a^{k+1}|f(b)|}{\varphi(a^{k+1}|f(b)|)}\bigg)^{s}\leq\Big(c\,\frac{a}{\varphi(a)}\log k\Big)^{s}\eta\log x.$ (1.5)
@@
$\sum_{\begin{subarray}{c}|b|\leq\eta\log x\\ b\,\neq\,b_{1},\ldots,b_{k}\end{subarray}}\bigg(\frac{a^{k+1}|f(b)|}{\varphi(a^{k+1}|f(b)|)}\bigg)^{s}\leq\Big(c\,\frac{a}{\varphi(a)}\log s\Big)^{s}\eta\log x.$ (1.6)
@@
Theorem 1.4 extends a result of Maynard ([3, Lemma 8.1]) which showed the same result but with $s=1$ and extends Theorem 1.4 in [4] which showed the inequality (1.6) but with an upper bound
@@
$\Big\{p\leq x:\frac{f(p)}{\varphi(f(p))}>t\Big\}\leq c_{1}\textup{exp}(-\textup{exp}(c_{2}t))\pi(x).$ (1.8)
In particular, we can take f(x)=x1f(x)=x-1 or $f(x)=x^{2}+1$ in Theorem 1.5.
@@ 2. Notation
We reserve the letter $p$ for primes. In particular, the sum $\sum_{p\leq K}$ should be interpreted as being over all prime numbers not exceeding $K$. We denote the number of primes not exceeding $x$ by $\pi(x)$ and the number of primes not exceeding $x$ that are congruent to $a$ modulo $k$ by $\pi(x;k,a)$. Let #A\#A denote the number of elements of a finite set $A$. We write $\mathbb{Z}$ for the set of all integers, $\mathbb{N}$ for the set of all positive integers, and $\mathbb{P}$ for the set of all primes. Let $(a_{1},\ldots,a_{n})$ be the greatest common divisor of integers $a_{1},\ldots,a_{n}$ and $a_{1},\ldots,a_{n}]$ their least common multiple. For real numbers $x,$ $y$ we also use $(x,y)$ to denote the open interval and $[x,y]$ to denote the closed interval. The usage of the notation will be clear from the context.
@@
$\frac{n}{\varphi(n)}\leq\frac{c}{\alpha}\prod_{p|n:\,\,p\leq(\log n)^{\alpha}}\biggl(1+\frac{1}{p}\biggr).$
@@
$\frac{a_{n}}{\varphi(a_{n})}$ $\leq\frac{c}{\alpha}\prod_{p|a_{n}:\,\,p\leq(\log a_{n})^{\alpha}}\biggl(1+\frac{1}{p}\biggr)$
@@
$=\Big(\frac{c}{\alpha}\Big)^{s}\sum_{d_{1},\ldots,d_{s}\in\mathcal{D}}\frac{\mu^{2}(d_{1})\ldots\mu^{2}(d_{s})}{d_{1}\ldots d_{s}}\sum_{\begin{subarray}{c}1\leq n\leq N:\\ d_{1}|a_{n},\ldots,\,d_{s}|a_{n}\end{subarray}}1$
@@
It is easy to see that $f(n)$ is supported on $\mathcal{D}$. Suppose that $m,n\in\mathcal{D}$ and $(m,n)=1$. For any $d_{1},\ldots,d_{s}\in\mathcal{D}$ such that $d_{1},\ldots,d_{s}]=mn$ there are unique $d_{i}^{{}^{\prime}},d_{i}^{{}^{\prime\prime}}\in\mathcal{D},$ $i=1,\ldots,s,$ such that $d_{i}^{{}^{\prime}}d_{i}^{{}^{\prime\prime}}=d_{i}$, $(d_{i}^{{}^{\prime}},d_{i}^{{}^{\prime\prime}})=1$, $d_{1}^{{}^{\prime}},\ldots,d_{s}^{{}^{\prime}}]=n$, and $d_{1}^{{}^{\prime\prime}},\ldots,d_{s}^{{}^{\prime\prime}}]=m$. Hence
@@
Since $\omega(n)\leq Kg(n)$ for any $n\in\mathcal{D}$, we have
$\sum_{n=1}^{N}\Big(\frac{a_{n}}{\varphi(a_{n})}\Big)^{s}$ $\leq K\Big(\frac{c}{\alpha}\Big)^{s}\sum_{n\in\mathcal{D}}f(n)g(n)=K\Big(\frac{c}{\alpha}\Big)^{s}\prod_{p\leq y}\Big(1+f(p)g(p)\Big)$
$=K\Big(\frac{c}{\alpha}\Big)^{s}\prod_{p\leq y}\Big(1+((1+p^{-1})^{s}-1)g(p)\Big).$
@@
$\sum_{n=1}^{N}\bigg(\frac{a_{n}}{\varphi(a_{n})}\bigg)^{s}$ $\leq\gamma N\Big(\frac{c}{\alpha}\Big)^{s}\prod_{p\leq y}\Big(1+((1+p^{-1})^{s}-1)g(p)\Big)$
$\leq\gamma N\Big(\frac{c}{\alpha}\Big)^{s}\prod_{p}\Big(1+((1+p^{-1})^{s}-1)g(p)\Big).$ (3.1)
@@
$\prod_{p>s}\Big(1+((1+p^{-1})^{s}-1)g(p)\Big)$ $\leq\prod_{p>s}\Big(1+\frac{esg(p)}{p}\Big)$
$\leq\textup{exp}\Big(es\sum_{p>s}\frac{g(p)}{p}\Big)\leq\textup{exp}(eLs).$ (3.3)
@@
$1+((1+p^{-1})^{s}-1)g(p)<1+c_{0}(1+p^{-1})^{s}\leq(1+c_{0})(1+p^{-1})^{s}\leq(1+c_{0})\textup{exp}\Big(\frac{s}{p}\Big).$
@@
$\prod_{p\leq s}\Big(1+((1+p^{-1})^{s}-1)g(p)\Big)$ $\leq(1+c_{0})^{s}\textup{exp}\Big(\sum_{p\leq s}\frac{s}{p}\Big)$
$\leq(1+c_{0})^{s}\textup{exp}(s\log\log(s+2)+c_{3}s),$ (3.4)
where $c_{3}>0$ is an absolute constant. From (3.1), (3), and (3.4) we obtain
@@
$\Big\{n\leq N:\frac{a_{n}}{\varphi(a_{n})}>t\Big\}\leq\textup{exp}(s\log\log(s+2)+Cs-s\log t)N.$ (3.5)
@@
$\Big\{n\leq N:\frac{a_{n}}{\varphi(a_{n})}>t\Big\}\leq\textup{exp}(-\textup{exp}(c_{2}t))N,$
@@
$\#\Big\{n\leq N:\frac{a_{n}}{\varphi(a_{n})}>t\Big\}$ $\leq N\leq N\,\frac{\textup{exp}(\textup{exp}(c_{2}t_{0}))}{\textup{exp}(\textup{exp}(c_{2}t))}$
@@
$\Big\{n\leq N:\frac{a_{n}}{\varphi(a_{n})}>t\Big\}\leq c_{1}\textup{exp}(-\textup{exp}(c_{2}t))N$
@@
where $\gamma=2cd$. The function $g(n)=n^{-1/d}$ is multiplicative. Also, $g(p)=p^{-1/d}\leq 1$ for any prime $p$ (hence, we can take $c_{0}=1$) and
$L=\sum_{p}\frac{g(p)}{p}=\sum_{p}\frac{1}{p^{1+1/d}}<\infty.$
@@
$\Big\{n\leq x:\frac{f(n)}{\varphi(f(n))}>t\Big\}\leq\frac{\widetilde{c}_{1}}{\textup{exp}(\textup{exp}(\widetilde{c}_{2}t))}x.$
@@
$S=\sum_{b\in\Omega}\bigg(\frac{a^{k+1}|f(b)|}{\varphi(a^{k+1}|f(b)|)}\bigg)^{s}\leq\Big(\frac{a}{\varphi(a)}\Big)^{s}\sum_{b\in\Omega}\bigg(\frac{|f(b)|}{\varphi(|f(b)|)}\bigg)^{s}.$ (4.3)
@@
$g(p^{\beta})=\begin{cases}\min(p,k)/p,&\text{if $\beta=1$;}\\ 1,&\text{if $\beta\geq 2$};\end{cases}$
@@
$\sum_{b\in\Omega}\bigg(\frac{|f(b)|}{\varphi(|f(b)|)}\bigg)^{s}\leq c^{s}(\eta\log x)\prod_{p\leq y}\Big(1+((1+p^{-1})^{s}-1)g(p)\Big)$
$\leq c^{s}(\eta\log x)\prod_{p}\Big(1+((1+p^{-1})^{s}-1)g(p)\Big)$ (4.4)
@@
$S\leq\Big(c\,\frac{a}{\varphi(a)}\Big)^{s}\eta\log x\prod_{p}\Big(1+((1+p^{-1})^{s}-1)g(p)\Big).$
Suppose that $s\leq k$ (we recall that $2\leq k<\log\log x$). If $p>k$, then $p>s$ and (3.2) holds. Since $g(p)=k/p$, we obtain
$\prod_{p>k}\Big(1+((1+p^{-1})^{s}-1)g(p)\Big)\leq\prod_{p>k}\Big(1+\frac{eks}{p^{2}}\Big)\leq c_{1}^{s},$
@@
$\prod_{p\leq k}\Big(1+((1+p^{-1})^{s}-1)g(p)\Big)\leq\textup{exp}\Big(\sum_{p\leq k}\frac{s}{p}\Big)\leq(c_{2}\log k)^{s},$
@@
Suppose that $s>k$ and $2\leq k<\log\log x$. Since $g(p)\leq k/p<s/p$ for any prime $p$, we obtain
$\prod_{p>s}\Big(1+((1+p^{-1})^{s}-1)g(p)\Big)\leq\prod_{p>s}\Big(1+\frac{es^{2}}{p^{2}}\Big)\leq c_{3}^{s}.$
Since $g(p)\leq 1$ for any prime $p$, we have
$\prod_{p\leq s}\Big(1+((1+p^{-1})^{s}-1)g(p)\Big)\leq\textup{exp}\Big(\sum_{p\leq s}\frac{s}{p}\Big)\leq(c_{4}\log s)^{s}.$
@@
We set $a_{p}=f(p)$. It is clear that $N=\pi(x)$. There exists a positive constant $\tau$ depending only on $f$ such that $f(p)\leq\tau p^{d}$ for any prime $p$. We take $M=\tau x^{d}$ and $\alpha=1/2$. We assume that $x\geq x_{0}$, where $x_{0}>0$ is a large constant depending only on $f$. By (4.1) and (4.2) we have $y=(\log M)^{\alpha}\leq(\log x)^{3/4}$ and $n\leq\sqrt{x}$ for any $n\in\mathcal{D}$.
We need the following result.
@@
If $1\leq k<x$ and $(a,k)=1$, then
$\pi(x;k,a)<\frac{3x}{\varphi(k)\log(x/k)}.$
Proof.
This is [1, Theorem 3.8]. ∎
Fix $n\in\mathcal{D}$, and let $\alpha_{1},\ldots,\alpha_{l}$ be all solutions of the congruence $f(x)\equiv 0$ (mod $n$). We have
$l=\rho(f,n)=\prod_{p|n}\rho(f,p)=\prod_{\begin{subarray}{c}p|n\\ p\nmid b_{d}\end{subarray}}\rho(f,p)\cdot\prod_{\begin{subarray}{c}p|n\\ p|b_{d}\end{subarray}}\rho(f,p)=A\cdot B.$
For any prime $p\nmid b_{d}$, by Lagrange’s theorem we have $\rho(f,p)\leq\min(p,d)$. Trivially $\rho(f,p)\leq p$ for any prime $p$. Hence,
$A\leq\prod_{\begin{subarray}{c}p|n\\ p\nmid b_{d}\end{subarray}}\min(p,d)\leq\prod_{p|n}\min(p,d),\qquad B\leq\prod_{\begin{subarray}{c}p|n\\ p|b_{d}\end{subarray}}p\leq\prod_{p|b_{d}}p=c(f).$
We obtain
$l\leq c(f)\prod_{p|n}\min(p,d),$ (5.1)
where $c(f)>0$ is a constant depending only on $f$.
@@
Fix $i\in\{1,\ldots,l\}$ and suppose that $(\alpha_{i},n)=1$. We have
$\log\Big(\frac{x}{n}\Big)\geq\log(\sqrt{x})=\frac{1}{2}\log x.$
Applying Lemma 5.1, we obtain
$\pi(x;n,\alpha_{i})\leq\frac{6x}{\varphi(n)\log x}.$ (5.3)
@@
$\pi(x;n,\alpha_{i})\leq 1\leq\frac{6x}{\varphi(n)\log x},$
@@
$\frac{6x}{\varphi(n)\log x}\geq\frac{6x}{n\log x}\geq\frac{6\sqrt{x}}{\log x}>1,$
if $x_{0}$ is large enough. We see that the inequality (5.3) holds in both cases $(\alpha_{i},n)=1$ and $(\alpha_{i},n)>1$. From (5.1) – (5.3) we get
$\omega(n)\leq\frac{6xl}{\varphi(n)\log x}\leq\frac{\gamma\pi(x)}{\varphi(n)}\prod_{p|n}\min(p,d),$
where $\gamma=\gamma(f)>0$ is a constant depending only on $f$.
@@
$g(p^{\beta})=\begin{cases}\min(p,d)/(p-1),&\text{if $\beta=1$;}\\ 1,&\text{if $\beta\geq 2$};\end{cases}$
@@
$L=\sum_{p}\frac{g(p)}{p}=\sum_{p\leq d}\frac{1}{p-1}+\sum_{p>d}\frac{d}{p(p-1)}<\infty,$
by Theorem 1.2 there exist positive constants $C$, $c_{1}$ and $c_{2}$ depending only on $f$ such that the inequalities (1.7) and (1.8) hold.
@@
$\Big\{p\leq x:\frac{f(p)}{\varphi(f(p))}>t\Big\}\leq\frac{\widetilde{c}_{1}}{\textup{exp}(\textup{exp}(\widetilde{c}_{2}t))}\pi(x).$